# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

### v1

@@ Abstract
Video retrieval requires aligning visual content with corresponding natural language descriptions. In this paper, we introduce Modality Auxiliary Concepts for Video Retrieval (MAC-VR), a novel approach that leverages modality-specific tags – automatically extracted from foundation models – to enhance video retrieval. We propose to align modalities in a latent space, along with learning and aligning auxiliary latent concepts, derived from the features of a video and its corresponding caption. We introduce these auxiliary concepts to improve the alignment of visual and textual latent concepts, and so are able to distinguish concepts from one other.
@@
We conduct extensive experiments on five diverse datasets: MSR-VTT, DiDeMo, TGIF, Charades and YouCook2. The experimental results consistently demonstrate that modality-specific tags improve cross-modal alignment, outperforming current state-of-the-art methods across three datasets and performing comparably or better across the other two.
I Introduction
@@
The emergence of prominent video-sharing platforms like YouTube and TikTok has supported uploading of millions of videos daily. The demand for better video retrieval methods, which align textual queries with relevant video content, has subsequently increased. Most existing works use two main approaches. The first [1, 2, 3] exclusively uses word and frame features. However, the second [4, 5, 6, 7, 8, 9] uses additional multi-modal information from videos – such as audio, speech, and objects – that is encoded and used for feature aggregation. In real-world scenarios, online videos often come with related textual information, such as tags – keywords associated with a video that describe its content and make it easier to search/filter. Few works [10, 11, 12] extract and exploit tags in video retrieval to better align the visual and textual modalities. These works only focus on extracting tags from the visual modality using pre-trained experts or a predefined visual token vocabulary. Inspired by these, we develop a novel method called MAC-VR that integrates multi-modal information by independently extracting relevant tags for both videos and texts without any manual annotation, utilizing the extensive knowledge from pre-trained Vision-Language Models (VLM) and Large Language Models (LLM) as shown in Fig. 1. The example in Fig. 1 shows the query “a girl doing gymnastics in the front yard”, the extracted visual (VT) and textual (TT) tags include sports, physical, outdoors, and outside can help align this video to the corresponding caption.
@@
Figure 1: Tags are extracted from both videos by VLM and texts by LLM using custom prompts designed to generate the most relevant tags for each modality. For example, visual and textual tags like sports and physical can help align this video to the corresponding caption. We learn latent auxiliary concepts from these tags, that help align the videos and texts.
@@
We build on the recent work DiCoSA [3], due to its performance and efficiency, where the visual and textual coarse features are split into compact latent factors which explicitly encode visual and textual concepts. Our MAC-VR extends this work by introducing additional (i.e. auxiliary) modality-specific latent concepts. We learn these auxiliary concepts from visual and textual tags using modality-specific foundation models. These are aligned to the latent concepts directly extracted from video and text, through an introduced Alignment Loss.
@@
Recently, many works [13, 3, 14, 15, 16, 10] use different inference strategies to improve the final video retrieval performance such as Querybank Normalisation (QB) [17] and Dual Softmax (DSL) [18]. In this paper, we analyse the impact of such strategies on our MAC-VR architecture to ensure fair comparison with state-of-the-art (SOTA) methods. Our results show that auxiliary concepts of both modalities, in addition to the Alignment Loss, help boost the retrieval performance and better distinguish the latent concepts.
@@
Our contribution is as follows: (i) We propose automatic extraction of modality-specific tags from foundational models to augment the video/text modalities. (ii) We use the tags to learn auxiliary latent concepts in each modality to extract meaningful representations. (iii) We propose a new Alignment Loss to better align and distinguish these learnt latent concepts. (iv) We analyse the impact of different inference strategies, extensively and fairly comparing our proposal with SOTA methods. (v) We conduct experiments on five datasets: MSR-VTT, DiDeMo, TGIF, Charades, and YouCook2. Across all datasets, the addition of our auxiliary concepts improves performance over our baseline. A detailed ablation on MSR-VTT verifies our design.
Figure 2: Examples of visual and textual tags (middle) for videos (left) and corresponding captions (right). Blue and orange highlight meaningful tags.
II Related works
Video Retrieval. Video retrieval learns an embedding for video and text to establish effective connections between video content and natural language descriptions. Early approaches [4, 5, 6, 7, 8, 9, 19, 20, 21, 22] relied on pre-trained features and/or multi-modal information inherent in videos, such as audio or speech, specialized to bridge the gap between videos and text. Notably, MMT [5] explores multi-modal data extracted by seven pre-trained experts but integrates them without explicit guidance, employing a brute-force method. Input modalities have also been masked, e.g. in [7], where the method can learn robust representations that enhance cross-modal matching; or video embeddings, such as in [19], by local temporal context, i.e. neighbouring actions. On the contrary, MAC-VR uses only video and text modalities without considering any additional modalities, such as audio or speech, or context.
Recent advancements in video retrieval have followed two main methodologies. The first involves extensive pre-training of models on large-scale video-text datasets, [23, 24]. The second focuses on transferring knowledge from image-based CLIP models [25] trained on extensive image-text pairs [21, 1, 26, 2, 14, 27, 28, 29, 30, 31, 3, 32, 33]. Some works [28, 30] use a distillation approach where a large network is first trained as a teacher network and then a smaller network is trained as a student network. In contrast, MAC-VR does not use any distillation approach and is trained directly by introducing auxiliary modality-specific tags. Similar to [3], where learnable queries and latent concepts are learnt during training, we learn latent auxiliary concepts from our modality-specific tags in addition to visual and textual latent concepts and use them as additional features to align the visual and textual concepts.
Vision-Language and Large Language Models in Image and Video Retrieval. The integration of Vision-Language Models (VLM) [34, 35, 36, 37, 38] and Large Language Models (LLM) [39, 40, 41, 42] in image [43, 44, 45, 46, 47] and video retrieval [48, 49, 12, 50, 51, 52, 53] has enabled significant advancements, showcasing the impressive understanding capabilities of these models. In [49] the authors demonstrate that LLMs can enhance the understanding via generation of video content, transferring their rich semantic knowledge. In [48], the authors explore how additional captions can enhance video retrieval by providing richer semantic context and improving matching accuracy between textual queries and video content. In contrast to these works, we do not generate additional captions as in [49, 48] but we leverage pre-trained VLMs and LLMs to generate words (i.e. visual and textual tags) that highlight relevant aspects of the action shown in the video and described by the caption. Our ablations verify the effectiveness of tags over additional captions extracted from the same foundation models.
Tags in Image and Video Understanding. The notion of tags in image and video understanding has found application across various tasks, including Video Retrieval [11, 10, 12], Video Moment Retrieval [54, 55], Video Recognition [56, 57], Fashion Image Retrieval [58, 59, 60, 61, 62] and Image Retrieval [63, 64, 65, 66, 67]. In [11, 10, 12], only the visual modality is considered to extract tags. Differently, MAC-VR considers both visual and textual modality to extract tags. More precisely, [11, 10] use pre-trained experts to extract tags from various modalities of videos, including object, person, scene, motion, and audio. In contrast, MAC-VR does not use pre-trained expert models to extract tags from a video or any additional modality such as audio. However, we generate visual and textual tags directly from videos and captions by using VLM and LLM, respectively. Similar to us, [12] uses image-language models to translate the video content into frame captions, objects, attributes, and event phrases by using a pre-defined visual token vocabulary. MAC-VR does not generate any additional captions from frames, and does not have any pre-fixed visual token vocabulary, instead using both the raw video and caption input to extract tags automatically.
III Modality Auxiliary Concepts for Video Retrieval
@@
We first define cross-modal text-to-video retrieval and inference strategies in Sec. III-A before describing our tag extraction approach in Sec. III-B. Finally, in Sec. III-C, we introduce our MAC-VR architecture.

### v3

# Bristol, UK Leveraging Modality Tags for … Video Retrieval

@@ Abstract
Video retrieval requires aligning visual content with corresponding natural language descriptions. In this paper, we introduce Modality Auxiliary Concepts for Video Retrieval (MAC-VR), a novel approach that leverages modality-specific tags – automatically extracted from foundation models – to enhance video retrieval. We propose to align modalities in a latent space, along with learning and aligning auxiliary latent concepts derived from the features of a video and its corresponding caption. We introduce these auxiliary concepts to improve the alignment of visual and textual latent concepts, allowing concepts to be distinguished from one another.
@@
We conduct extensive experiments on six diverse datasets: two different splits of MSR-VTT, DiDeMo, TGIF, Charades and YouCook2. The experimental results consistently demonstrate that modality-specific tags improve cross-modal alignment, outperforming current state-of-the-art methods across three datasets and performing comparably or better across others. Project Webpage: https://adrianofragomeni.github.io/MAC-VR/
1 Introduction
@@
The emergence of prominent video-sharing platforms, e.g., YouTube and TikTok, has supported uploading of millions of videos daily. The demand for better video retrieval methods, which align textual queries with relevant video content, has subsequently increased. Most existing works use two main approaches. The first [Fang et al.(2021)Fang, Xiong, Xu, and Chen, Luo et al.(2022b)Luo, Ji, Zhong, Chen, Lei, Duan, and Li, Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] exclusively uses word and frame features. However, the second [Dzabraev et al.(2021)Dzabraev, Kalashnikov, Komkov, and Petiushko, Gabeur et al.(2020)Gabeur, Sun, Alahari, and Schmid, Wang et al.(2021)Wang, Zhu, and Yang, Gabeur et al.(2022)Gabeur, Nagrani, Sun, Alahari, and Schmid, Liu et al.(2021)Liu, Fan, Qian, Chen, Ding, and Wang, Croitoru et al.(2021)Croitoru, Bogolin, Leordeanu, Jin, Zisserman, Albanie, and Liu] uses additional multi-modal information from videos – such as audio, speech, and objects – that is encoded and used for feature aggregation. In real-world scenarios, online videos often come with related textual information, such as tags – keywords associated with a video that describe its content and make it easier to search/filter. Few works [Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan, Wang et al.(2022a)Wang, Ge, Cai, Yan, Lin, Shan, Qie, and Shou, Wang et al.(2022c)Wang, Li, Xu, Zhou, Lei, Lin, Wang, Yang, Zhu, Hoiem, Chang, Bansal, and Ji] extract and exploit tags in video retrieval to better align the visual and textual modalities. These works only focus on extracting tags from the visual modality using pre-trained experts or a predefined visual token vocabulary. Inspired by these, we develop a novel method called MAC-VR that integrates multi-modal information by independently extracting tags for both videos and texts without any manual annotation, utilising the extensive knowledge from pre-trained Vision-Language Models (VLM) and Large Language Models (LLM) as shown in Fig. 1.
@@
Figure 1: Tags are extracted from both videos by VLM and texts by LLM using prompts designed to generate tags for each modality. For example, sports and physical tags can help align this video and its caption.
@@
The example in Fig. 1 shows the query “a girl doing gymnastics in the front yard”, the extracted visual (VT) and textual (TT) tags include sports, physical, outdoors, and outside, which can help align this video to its caption. We build on previous work [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen], due to its performance and efficiency, where visual and textual coarse features are split into compact latent factors which encode visual/textual concepts. MAC-VR extends this work by introducing auxiliary modality-specific latent concepts, which are learnt from visual and textual tags using modality-specific foundation models. These are aligned to the latent concepts extracted from videos and texts, through a novel Alignment Loss.
@@
Recently, many works [Liu et al.(2022)Liu, Xiong, Xu, Cao, and Jin, Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen, Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen, Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen, Ibrahimi et al.(2023)Ibrahimi, Sun, Wang, Garg, Sanan, and Omar, Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan] use different inference strategies to improve the final video retrieval performance such as Querybank Normalisation (QB) [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] and Dual Softmax (DSL) [Cheng et al.(2021)Cheng, Lin, Wu, Yang, and Shen]. In this paper, we analyse the impact of such strategies on our MAC-VR architecture to ensure fair comparison with state-of-the-art (SOTA) methods. Our results show that auxiliary concepts of both modalities, in addition to the Alignment Loss, help boost the retrieval performance and better distinguish the latent concepts.
@@
Our contribution is as follows: (i) We propose automatic extraction of modality-specific tags from foundational models to augment video/text. (ii) We use tags to learn auxiliary latent concepts in each modality to extract meaningful representations. (iii) We propose a new Alignment Loss to better align and distinguish these learnt latent concepts. (iv) We analyse the impact of different inference strategies, fairly comparing our proposal with SOTA methods. (v) We conduct experiments on six datasets: two splits of MSR-VTT, DiDeMo, TGIF, Charades, and YouCook2. Across all datasets, using auxiliary concepts improves performance over our baseline. A detailed ablation on MSR-VTT verifies our design.
2 Related works
Text-Video Retrieval. Video retrieval aligns video and text through learned embeddings. Early methods [Dzabraev et al.(2021)Dzabraev, Kalashnikov, Komkov, and Petiushko, Gabeur et al.(2020)Gabeur, Sun, Alahari, and Schmid, Wang et al.(2021)Wang, Zhu, and Yang, Gabeur et al.(2022)Gabeur, Nagrani, Sun, Alahari, and Schmid, Liu et al.(2021)Liu, Fan, Qian, Chen, Ding, and Wang, Croitoru et al.(2021)Croitoru, Bogolin, Leordeanu, Jin, Zisserman, Albanie, and Liu, Fragomeni et al.(2022)Fragomeni, Wray, and Damen, Zolfaghari et al.(2021)Zolfaghari, Zhu, Gehler, and Brox, Kunitsyn et al.(2022)Kunitsyn, Kalashnikov, Dzabraev, and Ivaniuta, Dong et al.(2022a)Dong, Li, Xu, Yang, Yang, Wang, and Wang] leveraged pre-trained features and multi-modal cues (e.g. audio, speech) to bridge the gap between videos and texts. Notably, MMT [Gabeur et al.(2020)Gabeur, Sun, Alahari, and Schmid] uses seven pre-trained experts in a brute-force fusion approach without explicit guidance. Input modalities have also been masked [Gabeur et al.(2022)Gabeur, Nagrani, Sun, Alahari, and Schmid] or enhanced by local temporal context [Fragomeni et al.(2022)Fragomeni, Wray, and Damen]. On the contrary, MAC-VR relies solely on video and text, without auxiliary modalities or context. Recent works have followed two main methodologies: involving extensive pre-training of models on large-scale video-text datasets, [Ge et al.(2022)Ge, Ge, Liu, Li, Shan, Qie, and Luo, Bain et al.(2021)Bain, Nagrani, Varol, and Zisserman] or transferring knowledge from image-based CLIP models [Radford et al.(2021a)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] trained on extensive image-text pairs [Kunitsyn et al.(2022)Kunitsyn, Kalashnikov, Dzabraev, and Ivaniuta, Fang et al.(2021)Fang, Xiong, Xu, and Chen, Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu, Luo et al.(2022b)Luo, Ji, Zhong, Chen, Lei, Duan, and Li, Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen, Huang et al.(2023)Huang, Gong, Pan, Jiang, Lv, Li, and Wang, Dong et al.(2023)Dong, Zhang, Zhang, Chen, Liu, Qu, Wang, and Liu, Guan et al.(2023)Guan, Pei, Shao, Liu, Li, Gu, Xu, Xu, Yan, and Lam, Tian et al.(2024)Tian, Zhao, Xin, Lan, and Li, Jin et al.(2024)Jin, Zhang, Gong, Xu, Deng, Wang, Zhang, Shen, and Feng, Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen, Xue et al.(2023)Xue, Sun, Liu, Fu, Song, Li, and Luo, Fang et al.(2023)Fang, Wu, Liu, Zhou, Song, Wang, Shu, Ji, and Wang]. Some employ knowledge distillation from large teacher networks [Dong et al.(2023)Dong, Zhang, Zhang, Chen, Liu, Qu, Wang, and Liu, Tian et al.(2024)Tian, Zhao, Xin, Lan, and Li], whereas MAC-VR bypasses distillation, training directly with auxiliary modality-specific tags. Similar to [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen], which learns latent queries and concepts, MAC-VR learns latent auxiliary concepts from these tags to enhance alignment between visual and textual features.
Vision-Language and Large Language Models in Image and Video Retrieval. The integration of Vision-Language Models (VLM) [Li et al.(2023b)Li, Li, Savarese, and Hoi, Liu et al.(2023a)Liu, Li, Wu, and Lee, Zhang et al.(2023b)Zhang, Li, and Bing, Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al., Zhang et al.(2023a)Zhang, Li, and Bing] and Large Language Models (LLM) [Touvron et al.(2023a)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al., Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al., Chiang et al.(2023)Chiang, Li, Lin, Sheng, Wu, Zhang, Zheng, Zhuang, Zhuang, Gonzalez, Stoica, and Xing, Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] has significantly advanced image [Qu et al.(2024)Qu, Li, Wang, Wang, Li, Nie, and Chua, Levy et al.(2023)Levy, Ben-Ari, Darshan, and Lischinski, Wang et al.(2024c)Wang, Elfardy, Dreyer, Small, and Bansal, Zhu et al.(2024)Zhu, Huang, Rudinac, and Kanoulas, Yan et al.(2023)Yan, Wang, Zhong, Dong, He, Lu, Wang, Shang, and McAuley] and video retrieval [Wu et al.(2023a)Wu, Luo, Fang, Wang, and Ouyang, Zhao et al.(2023)Zhao, Misra, Krähenbühl, and Girdhar, Wang et al.(2022c)Wang, Li, Xu, Zhou, Lei, Lin, Wang, Yang, Zhu, Hoiem, Chang, Bansal, and Ji, Shvetsova et al.(2024)Shvetsova, Kukleva, Hong, Rupprecht, Schiele, and Kuehne, Xu et al.(2024)Xu, Huang, Hou, Chen, Zhang, Feng, and Xie, Zhao et al.(2024)Zhao, Zhao, Zhou, Wu, Chu, Miao, Schroff, Adam, Liu, Gong, et al., Ventura et al.(2024)Ventura, Yang, Schmid, and Varol], showcasing their impressive understanding capabilities. Prior work enhances video retrieval by generating captions to enrich semantic context [Zhao et al.(2023)Zhao, Misra, Krähenbühl, and Girdhar, Wu et al.(2023a)Wu, Luo, Fang, Wang, and Ouyang]. In contrast, we leverage pre-trained VLMs and LLMs to generate visual and textual tags that highlight key aspects of the video and its caption. Ablation studies confirm the superiority of tags over new captions extracted from the same models (see Sec. 4.3).
Tags in Image and Video Understanding. Tags have been used across various tasks, including Video Retrieval [Wang et al.(2022a)Wang, Ge, Cai, Yan, Lin, Shan, Qie, and Shou, Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan, Wang et al.(2022c)Wang, Li, Xu, Zhou, Lei, Lin, Wang, Yang, Zhu, Hoiem, Chang, Bansal, and Ji], Video Moment Retrieval [Gao and Xu(2022), Wang et al.(2022b)Wang, Wu, Liu, and Yan], Video Recognition [Wu et al.(2023b)Wu, Wang, Luo, Wang, Yang, and Ouyang, Kahatapitiya et al.(2024)Kahatapitiya, Arnab, Nagrani, and Ryoo], Fashion Image Retrieval [Naka et al.(2022)Naka, Katsurai, Yanagi, and Goto, Wang et al.(2023a)Wang, Wang, Li, Li, Chen, Jin, Huang, Xiao, and Gao, Tian et al.(2023)Tian, Newsam, and Boakye, Shimizu et al.(2023)Shimizu, Nakamura, and Goto, Wahed et al.(2024)Wahed, Zhou, Yu, and Lourentzou] and Image Retrieval [Huang et al.(2024)Huang, Zhang, Ma, Tian, Feng, Zhang, Li, Guo, and Zhang, Liu et al.(2023b)Liu, Zheng, Wu, Tong, Liu, Chen, Wang, and Shen, Chaudhary et al.(2020)Chaudhary, Goyal, Goyal, and Chen, Zhu et al.(2021)Zhu, Cui, Cheng, Li, and Zhang, Chiquier et al.(2024)Chiquier, Mall, and Vondrick]. In [Wang et al.(2022a)Wang, Ge, Cai, Yan, Lin, Shan, Qie, and Shou, Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan, Wang et al.(2022c)Wang, Li, Xu, Zhou, Lei, Lin, Wang, Yang, Zhu, Hoiem, Chang, Bansal, and Ji], tags are extracted using pre-trained experts from various modalities of videos, including object, person, scene, motion, and audio. Differently, MAC-VR leverages both visual and textual inputs, generating tags directly from videos and captions using VLMs and LLMs without relying on expert models or additional modalities. While [Wang et al.(2022c)Wang, Li, Xu, Zhou, Lei, Lin, Wang, Yang, Zhu, Hoiem, Chang, Bansal, and Ji] uses image-language models with a fixed visual token vocabulary to derive frame-level captions and attributes, MAC-VR avoids predefined vocabularies or caption generation, instead extracting tags end-to-end from raw inputs.
3 Modality Auxiliary Concepts for Video Retrieval
@@
We first define text-to-video retrieval and inference strategies in Sec. 3.1 before describing our tag extraction approach in Sec. 3.2. Finally, in Sec. 3.3, we introduce MAC-VR.
3.1 Cross-Modal Text-to-Video Retrieval
Given a pair $(v_{i},t_{i})$, where $v_{i}$ represents a video and $t_{i}$ denotes its corresponding caption, the objective of Cross-Modal Text-to-Video Retrieval is to retrieve $v_{i}$ from a large gallery of videos, given the caption $t_{i}$ as query. Typically, models use two projection functions: $f_{v}:v_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ and $f_{t}:t_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ to map the respective modalities into a shared d-dimensional latent embedding space, $\Omega$. The aim is to align the representations in this space so that the representation of a video is close to that of its corresponding caption.
Figure 2: Examples of visual and textual tags (middle) for videos (left) and their captions (right). Blue and orange highlight semantic overlaps.
Following training, the standard inference strategy (IS) embeds a test video gallery and ranks these based on their distance from each query caption. Recent approaches utilise additional inference strategies to improve performance: Querybank Normalisation (QB) [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] and Dual Softmax (DSL) [Cheng et al.(2021)Cheng, Lin, Wu, Yang, and Shen] are the most popular. We introduce these strategies in Sec. B in supp.
3.2 Tag Extraction
We propose to extract tags automatically from either the video $v_{i}$, using a VLM, or the text in the caption $t_{i}$, using an LLM. These tags are word-level representations of common objects, actions, or general ideas present in the video or the caption. They can add additional useful information to retrieve the correct video given a text query as shown in Fig. 2. For instance, given the caption: a commercial for the Mazda 3 the car sliding around a corner, the general tags estimated from this caption are: product showcase, advertising technique, automotive marketing, which reflect the commercial. These words are terms that go beyond the exact caption but can help the model to understand the specific characteristics of this caption better.